OpenStudy (afloridagirl):

can someone check my answer?!?! Jane bought a car for $30,303. The value of the car depreciated at a constant rate per year. The table below shows the value of the car after the first and second years: Year 1 2 Value (in dollars) 25,757.55 21,893.92 Which function best represents the value of the car after t years? f(t) = 25,757.55(0.85)t f(t) = 30,303(0.85)t f(t) = 30,303 (0.15)t f(t) = 25,757.55 (0.15)t

OpenStudy (afloridagirl):

i think it is b

OpenStudy (afloridagirl):

but i think also c

OpenStudy (afloridagirl):

anyone?

OpenStudy (dallascowboys88):

me

OpenStudy (dallascowboys88):

ill check themfor you

OpenStudy (dallascowboys88):

it's b

OpenStudy (afloridagirl):

THANK YOU!

OpenStudy (dallascowboys88):

you were right the first time

OpenStudy (dallascowboys88):

No Prob:p;)

OpenStudy (dallascowboys88):

MEDAL!FAN!:P;)PLEASE?

OpenStudy (dallascowboys88):

THANK YOU SO MUCH!!!!!!!!!:P;)

OpenStudy (afloridagirl):

i love the dallas cowboys just saying

OpenStudy (afloridagirl):

@DallasCowboys88

OpenStudy (dallascowboys88):

Same Here HeHe:p;)

OpenStudy (afloridagirl):

can i ask another question really quick @DallasCowboys88

OpenStudy (afloridagirl):

please

OpenStudy (dallascowboys88):

Sure Why Not:p;)

OpenStudy (afloridagirl):

Given the two expressions shown below: square root of 4 plus square root of 25 square root of 4 plus square root of 2 Which statement best describes the two expressions? Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (afloridagirl):

i think it is a

OpenStudy (aliqanber):

A is rational, but B is irrational.

OpenStudy (aliqanber):

square root of 2 is irrational

OpenStudy (afloridagirl):

okay thank you can you explain how you got that

OpenStudy (dallascowboys88):

Correct See You Don't Even Need My Help Your Smart Without Me:p;)

OpenStudy (dallascowboys88):

afloridagirl@

OpenStudy (afloridagirl):

wait was it a? or was it c? @DallasCowboys88

OpenStudy (aliqanber):

A is just 2+5 = 7 --- rational B is 2 + sqrt(2) --- irrational since sqrt(2) is irrational (can't write it as a fraction)

OpenStudy (afloridagirl):

so c

OpenStudy (afloridagirl):

right?

OpenStudy (aliqanber):

100%

OpenStudy (afloridagirl):

okay thank you!

OpenStudy (afloridagirl):

What is 3 to the power of 3 over 2 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (afloridagirl):

if anyone could help with this

OpenStudy (afloridagirl):

i wanna say d

OpenStudy (afloridagirl):

@DallasCowboys88

OpenStudy (afloridagirl):

@aliqanber

OpenStudy (afloridagirl):

wait no its a

OpenStudy (aliqanber):

square root of 9

OpenStudy (afloridagirl):

so a?

OpenStudy (aliqanber):

b

OpenStudy (afloridagirl):

oh okay thank you

OpenStudy (aliqanber):

since: \[3^{\frac{ 3 }{ 2 }} = (3^{\frac{ 1 }{ 2 }})^{3} = \sqrt{3}^{3} = \sqrt{3^{3}} = \sqrt{27}\] sorry my bad, answer is d

OpenStudy (dallascowboys88):

💀💀💀💀💀💀💀💀💀💀💀

OpenStudy (dallascowboys88):

😎😎